# Using Half-Precision for GNN Training

### v1

@@ 5.1. GNN Training
This sub-section evaluates training results, showing accuracy, runtime, and memory usage. We focus on GCN, GAT, and GIN models to show the overall effectiveness of HalfGNN. These models contain intermediate feature lengths of 64. The intermediate feature length is important as the sparse kernels usually have this feature length during computation.
@@ 5.1.1. Training Accuracy
As we highlighted earlier in Fig. 1(c), the accuracy of DGL for half-precision is much lower than the float-based models, which questioned whether half-precision is a viable option for GNN training. Hence, we use DGL single-precision as the baseline in Fig. 4 to show that HalfGNN achieves similar accuracy to the baseline. The difference in accuracy is less than 0.3% for all except pubmed in GIN for which it is less than 1.0% improvement. Firstly, these differences are mostly due to random weight initialization, and changes by this margin if re-run. Secondly, we have made a minor adjustment to protect the additional overflow as discussed in section 4.3.1. This implies that half-precision is a viable option in place of float without compromising the GNN accuracy.
The reason for the similar accuracy in HalfGNN as float-based DGL is protecting the overflow through the proposed discretized reduction scaling in SpMM. Protection for SpMM is possible only at the sparse kernel level as frameworks themselves cannot protect against such overflow. To confirm this result, we replaced the discretized reduction with the usual reduction so that overflow would not be protected. This system did result in DGL-half like abnormal accuracy.
@@ 5.1.2. Training Runtime
Fig. 6 shows the runtime speedup of training time of GCN, GAT, and GIN using HalfGNN over DGL. All GNNs were trained for 400 epochs. Kindly recall that Fig. 1(a) and Fig. 1(b) have shown that half-precision based DGL kernels are slower than float-based DGL, while showing abnormal accuracy. Hence, we picked the faster (and correctly executed) baseline here. Compared to DGL float HalfGNN achieves 1.85×,3.55×1.85$\times$,3.55$\times$1.85 × , 3.55 ×, and 1.78×1.78$\times$1.78 × speedup on average for GCN, GAT, and GIN respectively, providing around an average of 2.30×2.30$\times$2.30 × speedup over float-based DGL
@@
Thus, the results so far show that half-based GNN is not only feasible(section 5.1.1) but is also significantly faster to train. The reason for the faster training time is the better sparse kernels of HalfGNN compared to prior GNN works, specifically introducing half-precision specific performance optimization to the sparse kernels. We show the runtime comparison of these kernels next by focusing on evaluating the impact of the proposed optimizations of HalfGNN.
5.1.3. Memory Consumption
HalfGNN also achieve 2.67$\times$ less memory consumption on average than float-based DGL considering all 3 GNNs as plotted in Fig. 5. While it is clear that half-precision would improve memory usage, one must note that the advantage is not purely due to moving to half-precision as DGL consumes excessive memory due to its design choices and has framework-memory overhead as shown recently (GNNBench, ).
@@ 5.2. Kernel Evaluation
In this section, we focus on individual kernel performances compared to established DGL/Cusparse kernel performance and compare with our results.
Figure 7. HalfGNN achieves 22.89×22.89$\times$22.89 × average speedup over DGL half-precision SpMM (Cusparse), and 7.12×7.12$\times$7.12 × average speedup over DGL half-precision SDDMM. Speedup over 64 has been clipped. (Log Scale, higher is better) )
@@ 5.2.1. SpMM Runtime
DGL uses Cusparse for SpMMve. We measure the runtime for feature sizes 32 and 64. We compared our runtime against the Cusparse half-precision. Fig. 7 shows the comparative speedup. We can observe that the speedup is massive for all datasets compared to the Cusparse. Furthermore, these speedups are consistent over feature sizes 32, and 64. HalfGNN SpMMve achieves over 64×64$\times$64 × runtime speedup in some cases, while achieving on average 22.89×22.89$\times$22.89 ×.
These speedups are observed due to two reasons: a) better design of half-precision kernel by HalfGNN which we have explored separately in section 5.3, and b) huge performance degradation by Cusparse half-precision library compared to its own float-based SpMM as discussed earlier in Fig. 1(a). Indeed, the huge speedup is due to the latter, as our additional measurement shows that HalfGNN achieves a more realistic 2.52×2.52$\times$2.52 × speedup over Cusparse-float.
SpMMv shares the same design principles and optimizations as SpMMve, whereas Cusparse does not provide this custom SpMM. Hence, we have not plotted it separately. This faster SpMM helps improve the training time for GCN, GAT, and GIN.
@@ 5.2.2. SDDMM Runtime
Fig. 7 shows that HalfGNN gets up to 12×12$\times$12 × speedup, while most datasets get over 6×6$\times$6 × speedup compared to the half-precision DGL SDDMM. The average speedup is around 7.12×7.12$\times$7.12 ×. The reason behind such massive speedup is that DGL half-precision kernels individually do not provide any runtime advantage over their float counterparts as it just replaces float data-type with half without even using half2 or any other optimizations, as shown earlier in Fig. 1(b). On the other hand, HalfGNN adopts half8 vector data-type which provides a huge performance boost in data-load time, increased instruction-level parallelism, and fewer rounds of inter-thread communication within a warp.
The speedup is huge for half-precision individual kernels of SpMM and SDDMM, yet we have shown only 2.30$\times$ speedup in training in Fig. 6. The realistic speedup is because we plotted float-based DGL in Fig. 6 as it trains faster than half-precision DGL, and trains correctly.
@@ 5.3. Micro-Benchmarking
In this section, we perform many micro-benchmarking to measure the impact of our proposed optimization choices.
@@ 5.3.1. SDDMM: Half8 vs Half2
We now show that usage of half8 is more efficient than using only half2 for SDDMM. Fig. 8 shows that using half8 can provide runtime speedup up to almost 3×3$\times$3 ×, and achieves an average speedup of 1.67×1.67$\times$1.67 × for feature sizes $32$ and $64$. Using half8 reduces the number of required inter-thread communications for the reduction in SDDMM and allows to issue load of four half2 features compared to just one , which results in better data-load performance due to increased instruction-level parallelism.
Figure 8. Using half8 achieves 1.67×1.67$\times$1.67 × speedup over half2 based SDDMM (Higher is better)
@@
Fig.  9 shows the runtime speedup achieved when HalfGNN removes the atomic write through intra-CTA communication, staging buffer, and follow-up kernel compared to a version that uses atomic write, keeping everything as is for edge-parallel design. This is true not only for Reddit and Ogb-product datasets but also for all the other datasets. Atomic implementations are more costly for half-precision than float data-type. A large part of our design for SpMMve revolved around how to get rid of atomic write. We believe, HalfGNN is the first-ever implementation, demonstrating the performance gain using non-atomic vectorized kernels in half-precision.
Figure 9. Speedup when atomic ops removed from HalfGNN SpMM
@@ 5.3.3. Huang-float vs Huang-Half2
. We chose Huang et al (huang2021understanding, ), a state-of-the-art workload-balanced vertex-parallel work, to showcase the effectiveness of our optimization proposed in HalfGNN. Huang et al SpMM groups the neighborhood of every vertex in the batch of 32, which is assigned to one warp, and hence uses atomic write to resolve conflicting writes in its float-version. We adopt it to half-precision, where we use half2, and remove the atomic write using staging buffer and follow-up kernel. However, the full application of HalfGNN SpMM does not apply to this approach specifically the edge-parallel approach. Further, a warp performs data loading edge features of 64 bytes and not 128 bytes due to batching only 32 neighbors thereby compromising the memory coalescing. Lastly, the data-load phase of edge features does not always start at even offset, leading to some compromise on the usage of half2 data type.
Figure 10. Huang-half2 gains 1.79×1.79$\times$1.79 × speedup over Huang-float.
Fig. 10 shows their runtime comparison showing that the adopted half-precision version achieves around 1.79×1.79$\times$1.79 × speedup on average. This is massive, as we achieve this without changing any fundamental nature of Huang et. al, without applying even the full range of proposed optimizations. This demonstrates the generality of our proposed optimization. Even though HalfGNN is an edge-parallel system, its design can massively improve performance for vertex-parallel systems also.
@@ 6. Conclusion
We studied GNN training in half-precision for the first time. We identified major issues that prevent existing half-precision systems from training GNN correctly. We provided generic solutions for that to explore half-precision training. Furthermore, we introduced novel techniques that can take advantage of half-precision to achieve state-of-the-art system performance for SpMM, SDDMM, and overall GNN training. We provided a system that integrates all these.

### v2

@@ 6.1. GNN Training
This sub-section evaluates training results, showing accuracy, runtime, and memory usage.
@@ 6.1.1. Training Accuracy
As we highlighted earlier in Fig. 1(c), the accuracy of DGL for half-precision is much lower than the float-based models. Hence, we use DGL single-precision as the baseline in Fig. 5 to show that HalfGNN achieves similar accuracy to the baseline. The improvement in accuracy is less than 0.3% for all except Pubmed in GIN for which it is less than 1.0% improvement. Firstly, half-precision works as a regularizer, resulting in minor accuracy improvements as discussed by mix-precision training (micikevicius2017mixed, ). Secondly, we have made a minor adjustment to protect the additional overflow in GIN as discussed in section 5.2.2. These accuracy results imply that half-precision is a viable option in GNN in place of the float if our proposed techniques are utilized.
Overflow Protection is the Key. Overflow protection through discretized reduction scaling in SpMM by HalfGNN is the key reason to achieve accuracy similar to DGL-float. To confirm this result, we replaced the discretized reduction with the usual reduction so that overflow would not be protected. This system did result in DGL-half like abnormal accuracy in GCN and GIN.
6.1.2. Memory Consumption
[…]
@@ 6.1.3. Training Runtime
[…]
Figure 7. HalfGNN training runtime speedup over DGL-half for feature size 64. (log scale, higher is better)
@@
HalfGNN speedup is greater compared to DGL-half than DGL-float. This confirms our earlier observation where Fig. 1(a) and Fig. 1(b) have shown that half-precision based DGL kernels are either slower (for SpMM) or have similar performance (for SDDMM) compared to the float-based DGL. Hence, by showing better performance over DGL-Float (faster and correctly executed baseline), HalfGNN truly advances the state-of-the-art.
The reason for the faster training time is due to the better sparse kernels of HalfGNN, specifically due to half-precision specific performance optimization to the sparse kernels as shown in section 6.2. Kindly note that GNN training involves many other kernels such as linear layer, activation, bias, softmax, etc., where baseline and HalfGNN use the same implementation of PyTorch. Hence, a sparse kernel speed-up (section 6.2), which is large for half-precision kernels of SpMM and SDDMM, still results in a significant average training time speed-up here.
@@ 6.2. Kernel Evaluation
This section focuses on individual kernel performances compared to established DGL/Cusparse kernels.
Figure 9. HalfGNN achieves $22.89\times$ average speedup over DGL half-precision SpMM (Cusparse), and $7.12\times$ average speedup over DGL half-precision SDDMM. Speedup over 64 has been clipped. (Log Scale, higher is better) )
@@ 6.2.1. SpMM Runtime
DGL uses Cusparse for SpMMve. We measure the runtime for feature sizes 32 and 64. We compared our runtime against the Cusparse half-precision. Fig. 9 shows the comparative speedup. We can observe that the speedup is massive for all datasets compared to Cusparse. Furthermore, these speedups are consistent over feature sizes 32 and 64. HalfGNN SpMMve achieves over $64\times$ runtime speedup in some cases, while achieving an average $22.89\times$ speedup.
These speedups are observed for two reasons. Firstly, Cusparse half-precision SpMM introduces huge performance degradation compared to its own float-based SpMM as discussed earlier in Fig. 1(a). Indeed, our additional measurement shows that HalfGNN achieves a more realistic $2.52\times$ average speedup over Cusparse-float (not plotted). Secondly, HalfGNN design (section 5.2) through its better data-load performance, increased computation throughput due to half2 usage, removal of atomic writes, workload balancing, etc., improves the HalfGNN kernel runtime.
Figure 10. HalfGNN SpMM achieves much higher SM and memory bandwidth utilization compared to DGL float and DGL half
To confirm, Fig. 10 plots a few performance counters using the Nvidia Nsight Compute (NCU) tool. HalfGNN achieves 80.92% average utilization of memory bandwidth compared to 20.22% in Cusparse-half (which is used by DGL-half) and 51.99% in Cusparse-float (which is used by DGL-float). The results show that HalfGNN has better data-load performance than DGL-half and DGL-float.
Compute or SM utilization for HalfGNN is 72.26% compared to 21.58% in Cusparse-half and 50.81% in Cusparse-float. While the improvement is significant, it shows that the SpMM kernel is not compute-bound, a well-known fact. These design choices are explored again using micro-benchmarks in section 6.3.
SpMMv shares the same design principles and optimizations as SpMMve, whereas Cusparse does not provide this custom SpMM. Hence, we have not plotted it separately.
@@ 6.2.2. SDDMM Runtime
Fig. 9 also shows that HalfGNN gets up to $12\times$ speedup, while most datasets get over $6\times$ speedup compared to the half-precision DGL SDDMM. The average speedup is around $7.12\times$. The reason behind such a large speedup is that DGL half-precision kernels individually do not provide any runtime advantage over their float counterparts, as it just replaces float data-type with half without even using half2 or any other optimizations, as shown earlier in Fig. 1(b). On the other hand, HalfGNN adopts half8 vector data-type which provides a huge performance boost in data-load time, and fewer rounds of inter-thread communication within a warp as discussed in section 5.1. Nvidia’s Compute Insight (NCU) tool confirms that HalfGNN achieves better memory bandwidth utilization of 83.71% compared to 50.85% in DGL-half and 50.59% in DGL-float as shown in Fig. 11.
Figure 11. HalfGNN SDDMM achieves an average of 83.71% memory bandwidth utilization compared to DGL float SDDMM (average of 50.59%), and DGL half SDDMM (average of 50.85%).
As DGL-half and DGL-float achieve similar memory bandwidth utilization, it empirically proves our hypothesis that merely replacing float with half would not provide any performance boost. Many of these design choices are evaluated separately in section 6.3.
@@ 6.3. Micro-Benchmarking
In this section, we perform many micro-benchmarks to measure the impact of a few proposed optimization choices.
@@ 6.3.1. SDDMM: Half8 vs Half2
We now show that usage of half8 is more efficient than using only half2 for SDDMM. Fig. 12 shows that using half8 can provide runtime speedup up to almost $3\times$, and achieves an average speedup of $1.67\times$ for feature sizes $32$ and $64$. Using half8 reduces the number of required inter-thread communications for the reduction in SDDMM and allows to issue load of four half2 features compared to just one. This results in better data-load performance due to increased instruction-level parallelism as discussed in section 5.1.
Figure 12. Using half8 achieves $1.67\times$ speedup over half2 based SDDMM (Higher is better)
@@
Figure 13. Speedup when atomic operations removed from HalfGNN SpMM
Fig. 13 shows the runtime speedup achieved when HalfGNN removes the atomic write through intra-CTA communication, staging buffer, and follow-up kernel compared to a version that uses atomic write, keeping everything as it is for edge-parallel design. This is true not only for Reddit and Ogb-product datasets, but also for all the other datasets. Atomic implementations are more costly for half-precision than float data-type. A large part of our design for SpMMve revolved around how to get rid of atomic write (section 5.2).
@@ 6.3.3. Huang-float vs Huang-Half2
We chose Huang et al. (huang2021understanding, ), a state-of-the-art workload-balanced vertex-parallel work, to showcase the generality and effectiveness of our optimization proposed in HalfGNN. We discussed the implementation ideas in section 5.4 except that we did not change the neighbor grouping size and kept it to its original 32. Hence, a warp performs data loading edge features of 64 bytes and not 128 bytes due to batching of 32 neighbors. So, the new prototype compromises the memory coalescing. This is not a fundamental limitation, as we can change its neighbor group size to 64 to overcome the issue.
Figure 14. Huang-half2 gains $1.79\times$ speedup over Huang-float for its SpMM design.
Fig. 14 shows their runtime comparison, showing that the adopted half-precision version achieves around $1.79\times$ speedup on average. This is significant, as we achieve this without changing any fundamental nature of Huang et al., without applying even the full range of proposed optimizations. This demonstrates the generality of our proposed optimization. Although HalfGNN is an edge-parallel system, its design can also equally improve the performance of vertex-parallel systems.
@@ 7. Conclusion
We studied GNN training in half-precision, where we identified major issues that prevent training GNN efficiently and correctly. We started with building a GNN system using the half2 vector data type by overcoming several challenges. We then provided new vector data types and discretized reduction in half-precision training to achieve state-of-the-art system performance for SpMM, SDDMM, and overall GNN training and accuracy. Furthermore, our work provided a system that integrates the proposed techniques and has shown that the proposed design techniques apply to both the edge-parallel and the vertex-parallel systems.
Acknowledgment
We would like to thank the anonymous reviewer for their effort and feedback. The work is supported in part by National Science Foundation grant 2245849. The views, opinions, and findings presented in this material are those of the authors and do not represent the official views or policies of the National Science Foundation.